OpenStudy (anonymous):

Does anyone know the how to solve this? which two equations are equivalent? A. y=(x +4) to the 2nd power y= x to the 2nd power + 16 B. y=(x -3) to the 2nd power y= x to the 2nd power -6x +9 C. y= (x +2) to the 2nd power y = x to the 2nd power + 2x +2 D. y=(x-6) to the 2nd power y= x to the 2nd power -12x - 36

OpenStudy (anonymous):

B because x-3)^2 = x^2-6x+9 lol

OpenStudy (anonymous):

nincompoop, Its been a while since i've seen these.. so I dont remember exactly..

OpenStudy (anonymous):

how did you get that or how did you came up with that jackson

OpenStudy (anonymous):

(X-3)(x-3) X*X=x^2 -3*x+(-3*x=-6x -3(-3)=9 x^2-6x+9

OpenStudy (anonymous):

and my names rodney

OpenStudy (anonymous):

sorry..

OpenStudy (anonymous):

okay, Thank you for your help =)

OpenStudy (anonymous):

sure